OpenStudy (anonymous):

A rectangle has sides with a length of 5 cm and 12 cm. What is the length of the diagonal of the rectangle? A. 17 cm B. 13 cm C. 12 cm D. 7 cm

OpenStudy (anonymous):

the Diagonal is just the third side of a triangle, so a^2 + b^2 = c^2 5^2 + 12^2 = c^2 25 + 144 = C^2 Square root of 169 = 13